OpenStudy (anonymous):

Given the following function, find: (a) vertex, (b) axis of symmetry, (c) intercepts, (d) domain, (e) range, (f) intervals where the function is increasing, (g) intervals where the function is decreasing f(x)=-7x^2-14x-3

OpenStudy (anonymous):

x=-b/2a x=-(-14)/2(-7) x=14/-14=-1 -7(-1)^2-14(-1)-3 -7+14-3=4 Vertex = (-1,4)

OpenStudy (anonymous):

axis of symetry will be -1